# Fine-Tuning Large Audio-Language Models with LoRA for Precise Temporal Localization of Prolonged Exposure Therapy Elements

### v2

@@ Abstract
Prolonged Exposure (PE) therapy is an effective treatment for post-traumatic stress disorder (PTSD), but evaluating therapist fidelity remains labor-intensive due to the need for manual review of session recordings. We present a method for the automatic temporal localization of key PE fidelity elements—identifying their start and stop times—directly from session audio and transcripts. Our approach fine-tunes a large pre-trained audio-language model, Qwen2-Audio, using Low-Rank Adaptation (LoRA) to process focused 30-second windows of audio-transcript input. Fidelity labels for three core protocol phases—therapist orientation (P1), imaginal exposure (P2), and post-imaginal processing (P3)—are generated via LLM-based prompting and verified by trained raters. The model is trained to predict normalized boundary offsets using soft supervision guided by task-specific prompts. On a dataset of 308 real PE sessions, our best configuration (LoRA rank 8, 30s windows) achieves a mean absolute error (MAE) of 5.3 seconds across tasks. We further analyze the effects of window size and LoRA rank, highlighting the importance of context granularity and model adaptation. This work introduces a scalable framework for fidelity tracking in PE therapy, with potential to support clinician training, supervision, and quality assurance.
Index Terms:
@@
Prolonged Exposure therapy, audio-language models, LoRA, QLoRA, temporal localization, multimodal learning, PTSD treatment, therapy fidelity, soft supervision.
I INTRODUCTION
@@
Post-Traumatic Stress Disorder (PTSD) is linked to multiple co-morbidities and can severely impact quality of life. Recovery is possible through treatments like Prolonged Exposure (PE) therapy, a gold-standard, evidence-based approach  [1]. However, trained PE clinicians are in short supply. Therapist fidelity—adherence to PE’s core components—is essential for effective outcomes [2], but assessing fidelity typically requires manual review of full sessions, making it time-consuming and hard to scale. While prior work has explored automated evaluation using speech and language processing [3, 4, 5, 6], most approaches focus on utterance-level classification rather than precise temporal localization of extended therapeutic phases—an essential requirement for PE fidelity assessment. Furthermore, the creation of large, finely annotated datasets remains a major bottleneck. This paper introduces a framework for second-level temporal localization of key PE therapy fidelity phases—specifically  [7, 8]: therapist orientation to imaginal exposure (P1), duration of imaginal exposure (P2), and post-imaginal processing (P3). Our approach leverages Qwen2 Audio [9], a model that jointly processes audio and transcript inputs, fine-tuned using Low-Rank Adaptation (LoRA) [10] and quantization techniques (QLoRA) [11] for computational efficiency. To address annotation challenges, we employ a soft-supervision pipeline in which initial timestamps are proposed by a large language model (LLM) using session transcripts and subsequently verified by trained raters. The model is trained on fixed-duration audio-transcript windows paired with task-specific prompts (e.g., “When did the Imaginal Exposure start?”) to predict the normalized offset of the target event. Accurate temporal resolution is therefore critical, as the PE manual specifies that the three major components of the session—the initial check-in (P1), the imaginal exposure (P2), and the processing of the imaginal exposure (P3)—should last approximately 10, 50, and 10 mins, respectively  [8, 7].
Our contributions are: (1) a method for fine-tuning Qwen2 Audio with LoRA for multimodal temporal localization in PE therapy; (2) a soft-supervision strategy that combines LLM-based annotation with trained rater verification; (3) an empirical analysis of how window duration and LoRA configurations affect localization accuracy; and (4) a demonstration of our method on real-world PE therapy data, achieving a mean absolute error (MAE) of 5.3s.
@@
To our knowledge, this is among the first efforts to fine-tune audio-language models for precise temporal localization of psychotherapy fidelity metrics. Section II reviews related work, Section III describes the dataset and annotation process, Section IV details our methodology, Section V presents experimental setup, Section VI has the results and discussion, and Section VII concludes.
II RELATED WORK
@@
Automated psychotherapy analysis enables scalable feedback, clinician training, and improved outcomes. We use pre-trained audio-language models with multimodal inputs to temporally localize fidelity metrics, using soft supervision with LLM-generated timestamps clinically verified for accuracy.
II-A Automated Evaluation of Psychotherapy
@@
II-B Temporal Event Localization in Audio and Speech
@@
II-C Large Pre-trained Audio-Language Models
@@
II-D Prompting and Few-Shot Learning in Multimodal Models
@@
II-E Annotation and Supervision Strategies
@@
II-F Research Contribution
While existing research has advanced psychotherapy analysis, multimodal learning, and temporal event detection, our work addresses temporal localization of psychotherapy fidelity metrics using pre-trained models, researcher-driven prompts, and soft-supervision—laying the foundation for fine-grained, scalable fidelity assessment in clinical contexts. Our work builds on recent advances in synthetic PE therapy datasets [32], fidelity evaluation frameworks [33], and empathy modeling in PTSD dialogue systems [34].
Figure 1: Timeline of a 70-minute PE therapy session showing typical segments for three fidelity metrics: P1 (orientation to imaginal exposure), P2 (imaginal exposure), and P3 (processing). Durations vary across sessions; this structure represents a common clinical pattern. All durations are in minutes.
III DATASET
We use a dataset of 318 PE therapy sessions (see Table II) recorded at Emory University, each lasting 60–90 minutes and capturing therapist-client interactions. All data were handled within IRB-approved environments, with no PII leaving secure storage. We annotated key protocol elements—orientation (P1), imaginal exposure (P2), and processing (P3)—which typically occur sequentially during a session (Figure 1), motivating our timestamp prediction task. To minimize manual labeling, we developed a zero-shot LLM pipeline that extracts protocol presence and timestamp boundaries from transcripts in JSON format. To ensure scalable validation, 10% (30 sessions) were reviewed by trained raters, yielding 94.4% timestamp accuracy (within 5–10 seconds) and 97.7% label accuracy. This validated our approach without requiring extensive expert clinician time. Audio was downsampled from 44.1–48 kHz to 16 kHz, normalized, and converted to .wav. Transcripts with sentence-level timestamps and speaker labels were generated using Amazon HealthScribe. After excluding sessions over 1.5 hours (due to transcription limits), those with alignment issues, or lacking valid labels for all protocol elements, we retained 308 recordings. The dataset was split into 216 training, 45 validation, and 47 test sessions, with balanced representation of therapy phases.
TABLE I: Example LLM-generated annotations with timestamps (in sec) and ratings for PE therapy phases.
ID Phase Description Start (s) Stop (s) Present
2 Imaginal lasted about 30–45 minutes (or about 15 for final imaginal). 105.83 2123.92 Yes
TABLE II: Summary statistics of the audio dataset. Durations are shown in HH:MM:SS format, with raw seconds in parentheses.
@@ III DATASET
Figure 2: Overview of our fidelity-aligned modeling pipeline. A task-specific prompt guides QLoRA fine-tuning on Qwen2 Audio using 30-120s audio-transcript windows randomly sampled around annotated start/stop points. The model is trained to predict normalized temporal offsets using clinically validated or synthetic labels.
TABLE III: Per-phase average, start, and end MAE (in seconds) for timestamp prediction on the test set. We run the experiments on three seeds (42, 78, 123) and report their Mean $\pm$ S.D.
@@
LoRA (r=4) $6.2\pm 1.3$ $6.2\pm 2.0$ $6.0\pm 1.2$ $5.4\pm 2.0$ $5.4\pm 2.2$ $5.5\pm 2.0$ $4.9\pm 1.5$ $5.2\pm 2.0$ $4.6\pm 1.4$
LoRA (r=8) $5.9\pm 1.4$ $5.5\pm 1.5$ $6.4\pm 1.4$ $5.0\pm 1.8$ $5.2\pm 1.8$ $4.9\pm 1.8$ $5.0\pm 0.5$ $5.5\pm 1.8$ $4.4\pm 0.8$
60s Head Only $12.2\pm 0.8$ $11.5\pm 1.0$ $12.9\pm 0.7$ $13.9\pm 0.5$ $13.4\pm 2.2$ $14.4\pm 1.2$ $13.7\pm 0.6$ $15.3\pm 1.1$ $12.2\pm 0.6$
LoRA (r=2) $11.3\pm 2.5$ $11.5\pm 2.9$ $11.2\pm 2.6$ $12.1\pm 3.1$ $12.8\pm 3.0$ $11.4\pm 3.3$ $12.1\pm 2.9$ $13.4\pm 2.6$ $10.8\pm 3.4$
LoRA (r=4) $11.9\pm 2.1$ $11.0\pm 2.1$ $12.7\pm 2.0$ $10.2\pm 2.0$ $10.5\pm 2.3$ $9.9\pm 2.1$ $11.6\pm 1.4$ $11.3\pm 1.0$ $11.8\pm 2.4$
@@
IV METHODOLOGY
IV-A Problem Formulation
[…]
The objective is to minimize the error between the predicted normalized offset $\hat{o}_{j}$ and the true normalized offset $o_{j}$.
IV-B Dataset and Soft-Supervision Labeling
Our processed dataset consists of 308 PE therapy sessions, preprocessed as described in Section III, and transcribed using Amazon HealthScribe, which provides sentence-level timestamps and speaker attribution. Ground-truth start and stop times for fidelity metrics P1, P2, and P3 were initially generated via soft supervision using Claude Sonnet 3.5, prompted to return timestamps in seconds as illustrated in Table I. These LLM-generated annotations were subsequently reviewed and corrected by trained trained raters to create verified ground-truth labels. All audio recordings were downsampled to 16kHz and split at the session level into train, validation, and test sets.
IV-C Model Architecture and Fine-tuning
@@
IV-D Training and Evaluation
The model minimizes MAE using the AdamW optimizer [35] (learning rate $1\times 10^{-4}$, weight decay 0.01, cosine schedule with 0.1 warmup ratio). Training runs for 10 epochs with a batch size of 1 and early stopping (patience = 3) based on validation MAE. We report MAE in seconds, calculated by denormalizing predicted offsets to absolute timestamps and comparing them to ground truth. Results are reported separately for each metric type (P1, P2, P3).
V EXPERIMENTAL SETUP
We implemented the temporal localization framework described in Section IV using Qwen2-Audio-7B-Instruct [9]. Experiments were conducted on a dataset of 308 PE therapy sessions, split into 216 train, 45 validation, and 47 test sessions as detailed in Section III. To assess the influence of context length and fine-tuning strategies, we varied (1) input audio-transcript window durations (30s, 60s, and 120s), and (2) LoRA ranks $r\in\{2,4,8\}$. These configurations were compared against a “Head Only” baseline where only the final regression layer was trained. All other QLoRA and training parameters (optimizer, learning rate, epochs, etc.) followed the specifications described in Section IV. Performance was evaluated using MAE (sec) for PE elements P1, P2, and P3.
VI RESULTS AND DISCUSSION
@@
Table III presents the mean $\pm$ standard deviation of MAE across three random seeds. Our best-performing configuration—30-second windows with LoRA rank $r=8$—achieved an average MAE of 5.3 seconds compared to ground truth across P1 ($5.9\pm 1.4$s), P2 ($5.0\pm 1.8$s), and P3 ($5.0\pm 0.5$s). This level of precision is promising for downstream clinical use, where timestamp-level fidelity tracking is essential. Earlier attempts at feature-level fusion using independent audio and text encoders produced overlapping embeddings across therapeutic phases, limiting their temporal resolution. In contrast, the end-to-end audio-language architecture of Qwen2-Audio-7B enabled learning of joint contextual representations that more effectively captured therapeutic cues. Shorter input windows (30s) consistently resulted in lower MAEs compared to longer windows. For instance, with the “Head Only” model, P1 Avg MAE increased from $6.8\pm 0.1$s (30s) to $12.2\pm 0.8$s (60s) and $25.2\pm 2.2$s (120s). This trend held across all configurations and phases, suggesting that temporally focused input helps the model better localize fidelity events. LoRA fine-tuning generally outperformed the “Head Only” baseline, especially at 60s and 120s. At 30s, LoRA $r=2$ and $r=8$ both reduced MAE for P2 Avg to around 5.1s and 5.0s, respectively, compared to 7.2s with “Head Only”. At 60s, LoRA $r=8$ consistently delivered the best results across phases (e.g., P1 Avg: $9.9\pm 0.1$s, P2 Avg: $9.7\pm 0.6$s, P3 Avg: $10.0\pm 0.6$s) and showed strong stability with low standard deviations. For 120s windows, the trend was less consistent. LoRA $r=2$ yielded the lowest P2 Avg MAE ($18.0\pm 2.3$s), outperforming both $r=4$ and $r=8$. Similarly, for P3, LoRA $r=4$ slightly outperformed $r=8$. These results suggest that higher LoRA ranks may overfit when the input spans become too broad, and smaller adapters can generalize better in such settings. The optimal LoRA rank varied depending on both window size and phase, highlighting the need for task-specific tuning. LoRA models also tended to show higher variability across seeds compared to “Head Only”, though configurations like LoRA $r=8$ at 60s achieved both low MAE and strong consistency. Overall, these results reinforce the benefits of combining temporally focused inputs with appropriately tuned model adaptation for timestamp prediction in therapy analysis.
VII CONCLUSION

### v4

@@ Abstract
Prolonged Exposure (PE) therapy is an effective treatment for post-traumatic stress disorder (PTSD), but evaluating therapist fidelity remains labor-intensive due to the need for manual review of session recordings. We present a method for the automatic temporal localization of key PE fidelity elements, identifying their start and stop times, directly from session audio and transcripts. Our approach fine-tunes a large pre-trained audio-language model, Qwen2-Audio, using Low-Rank Adaptation (LoRA) to process focused 30-second windows of audio-transcript input. Fidelity labels for three core protocol phases, therapist orientation (P1), imaginal exposure (P2), and post-imaginal processing (P3), are generated via LLM-based prompting and verified by trained raters. The model is trained to predict normalized boundary offsets using soft supervision guided by task-specific prompts. On a dataset of 308 real PE sessions, our best configuration (LoRA rank 8, 30s windows) achieves a mean absolute error (MAE) of 5.3s across tasks, within typical rater tolerance for timestamp review, enabling practical fidelity QC. We further analyze the effects of window size and LoRA rank, highlighting the importance of context granularity and model adaptation. This work introduces a privacy-preserving, scalable framework for fidelity tracking in PE therapy, with potential to support clinician training, supervision, and quality assurance.
@@
Index Terms—  Prolonged Exposure therapy, Temporal localization, Audio-language models, Multi-modal learning, Soft supervision
1 Introduction
@@
Post-Traumatic Stress Disorder (PTSD) is linked to multiple co-morbidities and can severely impact quality of life. Recovery is possible through treatments like Prolonged Exposure (PE) therapy, a gold-standard, evidence-based approach  [1]. However, trained PE clinicians are in short supply. Therapist fidelity, adherence to PE’s core components, is essential for effective outcomes [2], but assessing fidelity typically requires manual review of full sessions, making it time-consuming and hard to scale. While prior work has explored automated evaluation using speech and language processing [3, 4, 5, 6], most approaches focus on utterance-level classification rather than precise temporal localization of extended therapeutic phases, an essential requirement for PE fidelity assessment. Furthermore, the creation of large, finely annotated datasets remains a major bottleneck. This bottleneck effectively restricts high-quality care to a fraction of patients; thus, automating fidelity is not merely a technical convenience but an ethical imperative for healthcare equity. This paper introduces a framework for second-level temporal localization of key PE therapy fidelity phases, specifically [7, 8]: therapist orientation to imaginal exposure (P1), duration of imaginal exposure (P2), and post-imaginal processing (P3). Our approach leverages Qwen2 Audio [9], a model that jointly processes audio and transcript inputs, fine-tuned using Low-Rank Adaptation (LoRA) [10] and quantization techniques (QLoRA) [11] for computational efficiency. By distilling heavy server-side reasoning into a quantized local model, we ensure that patient audio, potentially the most sensitive biometric data, never leaves the secure inference environment, addressing a primary barrier to clinical AI adoption. To address annotation challenges, we employ a soft-supervision pipeline in which initial timestamps are proposed by a large language model (LLM) using session transcripts and subsequently verified by trained raters. We distill these high-quality annotations into a locally running model, avoiding external API calls and enabling privacy-preserving deployment in clinics or smartphones [12, 13]. The model is trained on fixed-duration audio-transcript windows paired with task-specific prompts (e.g., “When did the Imaginal Exposure start?”) to predict the normalized offset of the target event. Accurate temporal resolution is therefore critical, as the PE manual specifies that the three major components of the session, the initial check-in (P1), the imaginal exposure (P2), and the processing of the imaginal exposure (P3), should last approximately 10, 50, and 10 mins, respectively [7, 8]. By automating these specific timestamps, we enable a “human-in-the-loop” workflow where supervisors need only review flagged segments rather than full sessions, potentially reducing the fidelity assessment burden by orders of magnitude.
Our contributions are:
  1. 1.
    Proposed a method for fine-tuning Qwen2-Audio with LoRA for multimodal temporal localization in PE therapy.
  2. 2.
    Introduced a novel soft-supervision strategy that combines LLM-based annotation with rater verification.
  3. 3.
    Analyzed the impact of window duration and LoRA configuration choices.
  4. 4.
    Demonstrated the approach on real-world PE therapy data.
  5. 5.
    Presented a continuous temporal regression framework for fidelity tracking, building on advances in synthetic PE datasets [14], evaluation frameworks [15], and empathy modeling [16], providing a foundation for scalable, fine-grained clinical assessment.
@@
To our knowledge, this is among the first efforts to fine-tune audio-language models for precise temporal localization of psychotherapy fidelity metrics. The remainder of this paper is organized as follows: Section  2 reviews related work, Section 3 describes the dataset and annotation, Section 4 details our methodology, Section 5 presents experimental setup, Section 6 discusses the results, and Section 7 concludes.
2 RELATED WORK
@@
Automated psychotherapy analysis scales feedback, enhances clinician training, and improves outcomes. We fine-tune pre-trained audio-language models on multimodal inputs to temporally localize fidelity metrics, using soft-supervised LLM-generated timestamps verified for clinical accuracy.
2.1 Automated Evaluation of Psychotherapy
@@
2.2 Temporal Event Localization in Audio and Speech
@@
2.3 Large Pre-trained Audio-Language Models
@@
2.4 Multimodal Prompting & Few-Shot Learning
@@
2.5 Annotation and Supervision Strategies
@@
Fig. 1: Timeline of a 70-minute PE therapy session showing typical segments for three fidelity metrics: P1 (orientation to imaginal exposure), P2 (imaginal exposure), and P3 (processing). Durations vary across sessions; this structure represents a common clinical pattern. All durations are in minutes.
3 DATASET
We use 318 PE therapy sessions (Table 1) recorded at Emory University, 60-90 minutes each. All data were handled in IRB-approved environments; no PII left secure storage. We annotated three protocol phases: orientation (P1), imaginal exposure (P2), and processing (P3), which typically occur sequentially (Figure 1), motivating timestamp prediction. To reduce manual labeling, a zero-shot LLM extracted protocol presence and start/stop times from transcripts (JSON). For scalable validation, trained raters reviewed 10% (30 sessions), yielding 94.4% timestamp accuracy (within 5-10 s) and 97.7% label accuracy. Audio was downsampled from 44.1-48 kHz to 16 kHz, normalized, and saved as WAV; transcripts with sentence-level timestamps and speaker labels were generated using Amazon HealthScribe. After excluding sessions over 1.5 hours, sessions with alignment issues, or those lacking valid labels, 308 recordings remained. We split the data into 216/45/47 train/validation/test sessions with balanced phase coverage.
Table 1: Summary statistics of the audio dataset. Durations are shown in $H$:$M$:$S$ format, seconds in paranthesis.
@@ 3 DATASET
Table 2: Example LLM-generated annotations with timestamps (in sec) and ratings for PE therapy phases.
ID Phase Description Start (s) Stop (s) Present
2 Imaginal lasted about 30-45 minutes (or about 15 for final imaginal). 105.83 2123.92 Yes
Fig. 2: Overview of our fidelity-aligned modeling pipeline. A task-specific prompt guides QLoRA fine-tuning on Qwen2 Audio using 30-120s audio-transcript windows randomly sampled around annotated start/stop points. The model is trained to predict normalized temporal offsets using clinically validated or synthetic labels.
Table 3: Per-phase average, start, and end MAE (in seconds) for timestamp prediction on the test set. We run the experiments on three seeds (42, 78, 123) and report their Mean $\pm$ S.D.
@@
LoRA (r=4) $6.2\pm 1.3$ $6.2\pm 2.0$ $6.0\pm 1.2$ $5.4\pm 2.0$ $4\pm 2.2 5.5$$5\pm 2.0$ $4.9\pm 1.5$ $5.2\pm 2.0$ $4.6\pm 1.4$
LoRA (r=8) $9\pm 1.4 5.5$$5\pm 1.5$ $6.4\pm 1.4$ $5.0\pm 1.8$ $5.2\pm 1.8$ $4.9\pm 1.8$ $0\pm 0.5 5.5$$5\pm 1.8$ $4.4\pm 0.8$
60s Head Only $2.2\pm 0.8 11$$1.5\pm 1.0$ $12.9\pm 0.7$ $13.9\pm 0.5$ $13.4\pm 2.2$ $14.4\pm 1.2$ $13.7\pm 0.6$ $15.3\pm 1.1$ $12.2\pm 0.6$
LoRA (r=2) $1.3\pm 2.5 11$$5\pm 2.9 11.$$1.2\pm 2.6$ $12.1\pm 3.1$ $2.8\pm 3.0 11$$1.4\pm 3.3$ $12.1\pm 2.9$ $13.4\pm 2.6$ $10.8\pm 3.4$
LoRA (r=4) $1.9\pm 2.1 11$$1.0\pm 2.1$ $12.7\pm 2.0$ $10.2\pm 2.0$ $10.5\pm 2.3$ $9.9\pm 2.1$ $1.6\pm 1.4 11$$3\pm 1.0 11.$$1.8\pm 2.4$
@@
4 METHODOLOGY
4.1 Problem Formulation
We treat fidelity not as a sequence of discrete classifications, but as a continuous temporal regression task. This formulation allows the model to capture the fluid, non-binary transitions typical of therapeutic dialogue, which rigid segmentation often mischaracterizes. We formulate temporal localization of PE therapy fidelity metrics as a regression task. Given a PE therapy recording, transcript, and pre-defined metrics (P1: Orientation; P2: Imaginal exposure; P3: Post-imaginal processing), we predict precise start and end times of each metric instance. To do this, we process focused segments of the session. Let $W_{j}$ represent an input window $j$, comprising an audio segment $A_{j}$ and its corresponding transcript excerpt $T_{j}$. Each window $W_{j}$ is extracted around a known boundary (start or end) of a specific fidelity metric $q_{i}$. The true absolute time of this boundary within the full session is $t_{\text{abs}}$. The window $W_{j}$ has a duration $D_{j}$ and starts at an absolute time $t_{\text{start},j}$. The true boundary $t_{\text{abs}}$ is therefore located at a normalized offset $o_{j}=(t_{\text{abs}}-t_{\text{start},j})/D_{j}$ within this window, where $o_{j}\in[0,1]$. Our model $M$, a fine-tuned audio-language model, takes the audio segment $A_{j}$, transcript excerpt $T_{j}$, and a task-specific textual prompt $P(q_{i},\text{boundary\_type})$ (where boundary_type is start or end) as input. The model is trained to predict the normalized offset $\hat{o}_{j}=M(A_{j},T_{j},P(q_{i},\text{boundary\_type}))$. The objective is to minimize the error between the predicted normalized offset $\hat{o}_{j}$ and the true normalized offset $o_{j}$.
4.2 Dataset and Soft-Supervision Labeling
Our processed dataset consists of 308 PE therapy sessions, preprocessed as described in Section 3, and transcribed using Amazon HealthScribe, which provides sentence-level timestamps and speaker attribution. Ground-truth start and stop times for fidelity metrics P1, P2, and P3 were initially generated via soft supervision using Claude Sonnet 3.5, prompted to return timestamps in seconds as illustrated in Table 2. These LLM-generated annotations were subsequently reviewed and corrected by trained raters to create verified ground-truth labels. All downsampled audio recordings were split at the session level into train/validation/test sets.
4.3 Model Architecture and Fine-tuning
@@
5 EXPERIMENTAL SETUP
We implemented the framework described in Section 4 using Qwen2-Audio-7B-Instruct [9] on 308 PE sessions (Section 3). We also evaluated a “Head Only” baseline that trains only the final regression layer to isolate the impact of parameter-efficient fine-tuning on the joint multimodal embeddings. We prioritize this adaptation analysis over unimodal (text-only) ablation because purely text-based approaches are constrained by the discrete granularity of ASR timestamps, which typically lack the continuous resolution required for second-level precision in therapeutic phase transitions. Crucially, therapy phase transitions often occur in the silences or prosodic shifts between words, signals that text-only models inherently discard. Models were fine-tuned with QLoRA using the AdamW optimizer [37] (learning rate $1\times 10^{-4}$, weight decay 0.01, cosine schedule with 0.1 warmup ratio) for 10 epochs with a batch size of 1 and early stopping (patience=3) based on validation MAE. Predicted offsets were denormalized to absolute timestamps, and mean absolute error (MAE, in seconds) was reported for P1, P2, and P3.
6 RESULTS AND DISCUSSION
@@
Table 3 presents the mean $\pm$ standard deviation of MAE across three random seeds. Our best-performing configuration, 30-second windows with LoRA rank $r=8$, achieved an average MAE of 5.3 seconds compared to ground truth across P1 ($5.9\pm 1.4$s), P2 ($5.0\pm 1.8$s), and P3 ($5.0\pm 0.5$s). Notably, an MAE of 5.3s falls within the natural variance of human cognitive pauses. This suggests the model’s localization error is effectively indistinguishable from human rater variability for the purpose of clinical quality control. This level of precision is promising for downstream clinical use, where timestamp-level fidelity tracking is essential. Earlier attempts at feature-level fusion using independent audio and text encoders produced overlapping embeddings across therapeutic phases, limiting their temporal resolution. In contrast, the end-to-end audio-language architecture of Qwen2-Audio-7B enabled learning of joint contextual representations that more effectively captured therapeutic cues. Shorter input windows (30s) consistently resulted in lower MAEs compared to longer windows. For instance, with the “Head Only” model, P1 Avg MAE increased from $6.8\pm 0.1$s (30s) to $12.2\pm 0.8$s (60s) and $25.2\pm 2.2$s (120s). This reveals a critical context-granularity trade-off: while larger windows provide more semantic history, they dilute the temporal sharpness required for boundary detection. The 30s window emerges as the optimal balance, retaining sufficient rhetorical context while maximizing edge precision. LoRA fine-tuning generally outperformed the “Head Only” baseline, especially at 60s and 120s. At 30s, LoRA $r=2$ and $r=8$ both reduced MAE for P2 Avg to around 5.1s and 5.0s, respectively, compared to 7.2s with “Head Only”. At 60s, LoRA $r=8$ consistently delivered the best results across phases (e.g., P1 Avg: $9.9\pm 0.1$s, P2 Avg: $9.7\pm 0.6$s, P3 Avg: $10.0\pm 0.6$s) and showed strong stability with low standard deviations. For 120s windows, the trend was less consistent. LoRA $r=2$ yielded the lowest P2 Avg MAE ($18.0\pm 2.3$s), outperforming both $r=4$ and $r=8$. Similarly, for P3, LoRA $r=4$ slightly outperformed $r=8$. These results suggest that higher LoRA ranks may overfit when the input spans become too broad, and smaller adapters can generalize better in such settings. The optimal LoRA rank varied depending on both window size and phase, highlighting the need for task-specific tuning. LoRA models also tended to show higher variability across seeds compared to “Head Only”, though configurations like LoRA $r=8$ at 60s achieved both low MAE and strong consistency. Overall, these results reinforce the benefits of combining temporally focused inputs with appropriately tuned model adaptation for timestamp prediction in therapy analysis.
7 CONCLUSION